OpenStudy (anonymous):

What are the possible steps involved in solving the equation shown? Check all that apply: 3.5 + 1.2(6.3 - 7x) = 9.38 -Add 3.5 and 1.2. -Distribute 1.2 to 6.3 and -7x . -Combine 6.3 and -7x . -Combine 3.5 and 7.56. -Subtract 11.06 from both sides. -Divide both sides by 9.38.

OpenStudy (anonymous):

Distribute 1.2 to 6.3 and -7.x but im thinking that you can first add 3.5 and 1.2 but im again thinking that would be against pendas

OpenStudy (michele_laino):

I think Distribute 1.2 to 6.3 and –7x is the right option, furthermore there is another rigfht option

OpenStudy (michele_laino):

oops...there is another right option

OpenStudy (michele_laino):

hint: \[\begin{gathered} 3.5{\text{ }} + {\text{ }}1.2\left( {6.3{\text{ }}--{\text{ }}7x} \right){\text{ }} = {\text{ }}9.38 \hfill \\ 3.5{\text{ }} + {\text{ }}7.56 - 8.4x{\text{ }} = {\text{ }}9.38 \hfill \\ 11.06 - 8.4x = 9.38 \hfill \\ \end{gathered} \]

OpenStudy (anonymous):

Thanks for the help I got the answer now :)

OpenStudy (michele_laino):

Thanks!